OpenStudy (anonymous):

How many positive integers less than 100 have a remainder of 3 when divided by 7?

OpenStudy (anonymous):

14.

OpenStudy (anonymous):

\[3,7+3,2\times 7+3,3\times 7+3,...\]

OpenStudy (anonymous):

so abtrehearn is right. since \[14\times 7+3=101\] which tells us that there are 13 +1 because of the first one

OpenStudy (anonymous):

Are u counting zero? If so why?

OpenStudy (anonymous):

I'm thinking 13.

OpenStudy (anonymous):

because 7 goes in to 3 zero times with a remainder of 3

OpenStudy (anonymous):

I was thinking what LA-Mrc said but my book says the answer is 14. If 7 doesn't go into 3 wouldn't the remainder be 7 : ?

OpenStudy (anonymous):

Here's how I worked it out: 10 is the first positive integer that can be divide by 7 with a remainder of 3. The next is 17, then 24, ... all the way to 94.

OpenStudy (anonymous):

10, 17, 24, ..., 94 That should be 13 terms.

OpenStudy (anonymous):

The number 3 is also in the list of positive integers less than 100 whose remainder is 3 upon dividing by 7. The list is {3, 10, 17, ... , 94}.

OpenStudy (anonymous):

I guess I missed the first one, i.e., 3.

OpenStudy (anonymous):

Good job, abtrehearn,

OpenStudy (anonymous):

I'm not sure I understand how 3 works also : /

OpenStudy (anonymous):

Ten-four, IA-MrC :^)